# What Are They Talking About? A Benchmark of Knowledge-Grounded Discussion Summarization

### v2

@@ Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We gratefully acknowledge the research funding provided by Fanyu AI Laboratory. We thank the anonymous reviewers for their constructive feedback that helped improve this work.
@@ Appendix A Paradigm-Level Performance Visualization
Figure 6: Overall performance under the EBS-AOS paradigm. Models are stratified into three tiers: Tier-1 ([72,77][72,77][ 72 , 77 ]), Tier-2 ([64,68][64,68][ 64 , 68 ]), and Tier-3 ([34,46][34,46][ 34 , 46 ]).
@@ Appendix B Further Analysis
[…]
@@
[…]
@@ C.1 ABS Evaluation Metrics
Figure 8: Overall performance under the ABS-AOS paradigm. Models are stratified into three tiers: Tier-1 ([55,61][55,61][ 55 , 61 ]), Tier-2 ([41,49][41,49][ 41 , 49 ]), and Tier-3 ([25,27][25,27][ 25 , 27 ]).
Figure 9: Average overall performance across both paradigms. Models are stratified into three tiers: Tier-1 ([61,69][61,69][ 61 , 69 ]), Tier-2 ([53,56][53,56][ 53 , 56 ]), and Tier-3 ([29,37][29,37][ 29 , 37 ]).
@@ Appendix D Detailed Benchmark Statistics
[…]
@@
Among the 242824282428 atomic fact units decomposed from the 432432432 supporting paragraphs, 163816381638 are consistently annotated as key supporting atomic facts, 780780780 as non-key atomic facts, and 101010 as intra-repetitive atomic facts.
Among the 518751875187 atomic fact units from the 100510051005 nonsupporting paragraphs, 499649964996 are automatically annotated as nonsupporting atomic facts, 176176176 as masked conflicting facts, and 151515 as intra-repetitive atomic facts.
@@
A total of 873873873 clear atomic opinions are expert manually annotated. Of these, 800800800 contain clarified implicit references, while 737373 do not require clarification. Within the 800800800 opinions, there are a total of 111311131113 written clarified implicit references.
@@ Appendix E LLM Sources
Appendix F Prompts and Instructions
F.1 Unified Parameter Settings
In this work, all 12 evaluated LLMs use consistent parameter settings for prompts and instructions: max_token=4096 and temperature=0. No other default parameters are modified.
F.2 Atomic Fact Decomposition Instruction
@@ F.3 Structured Prompts
F.4 Self-Reflection Instructions
@@ F.5 Verification Prompts
F.6 Error Detection Instruction

### v3

@@ Acknowledgement
This research was supported in part by the National Natural Science Foundation of China (Grant Nos. 62276017, 62406033, U1636211, 61672081) and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLCCSE-2024ZX-18). We thank the anonymous reviewers for their feedback that helped improve this work.
@@ Appendix A Paradigm-Level Performance Visualization
Figure 6: Overall performance under the EBS-AOS paradigm. Models are stratified into three tiers: Tier-1 ($[72,77]$), Tier-2 ($[64,68]$), and Tier-3 ($[34,46]$).
@@ Appendix B Further Analysis
By analyzing the performance gap between the two summarization paradigms (see Figures 6, 7, and 8 in Appendix A), we find that different LLMs excel in different paradigms. For example, Claude 3.5 Sonnet exhibits a significant gap ($28.29$ $\%$ ), while Claude 3 Opus demonstrates a relatively smaller gap ($11.37$ $\%$ ), indicating stronger stability.
@@
As shown in Table 3, most LLMs demonstrate polarization in the increases and decreases between $\mathrm{SP}_{\mathrm{R}}$ and $\mathrm{SP}_{\mathrm{P}}$, as well as $\mathrm{KSAF}_{\mathrm{R}}$ and $\mathrm{KSAF}_{\mathrm{P}}$. This suggests different reflection strategies: some prioritize precision (e.g., DeepSeek-V3 with $\mathrm{SP}_{\mathrm{R}}$ $\downarrow 1.19$}, $\mathrm{SP}_{\mathrm{P}}$ $\uparrow 1.37$}), while others prioritize recall (e.g., Claude 3 Opus). A few models exhibit simultaneous increases (e.g., GPT-4o with $\mathrm{SP}_{\mathrm{R}}$ $\uparrow 1.56$}, $\mathrm{SP}_{\mathrm{P}}$ $\uparrow 0.34$}) or decreases (e.g., Qwen-Max), indicating more balanced or weaker reflection abilities.
Figure 8: Overall performance under the ABS-AOS paradigm. Models are stratified into three tiers: Tier-1 ($[55,61]$), Tier-2 ($[41,49]$), and Tier-3 ($[25,27]$).
Figure 9: Average overall performance across both paradigms. Models are stratified into three tiers: Tier-1 ($[61,69]$), Tier-2 ($[53,56]$), and Tier-3 ($[29,37]$).
@@ Appendix D Detailed Benchmark Statistics
Among the $1696$ paragraphs from $100$ original articles, the expert annotation consistency rate is $84.7$ $\%$ . A total of $1437$ paragraphs are retained for the final news articles, while $208$ are identified as ambiguous paragraphs and removed. Of the $1437$ retained, $432$ are annotated as supporting and $1005$ as nonsupporting.
@@
Among the $2428$ atomic fact units decomposed from the $432$ supporting paragraphs, $1638$ are consistently annotated as key supporting atomic facts, $780$ as non-key atomic facts, and $10$ as intra-repetitive atomic facts.
Among the $5187$ atomic fact units from the $1005$ nonsupporting paragraphs, $4996$ are automatically annotated as nonsupporting atomic facts, $176$ as masked conflicting facts, and $15$ as intra-repetitive atomic facts.
@@
A total of $873$ clear atomic opinions are expert manually annotated. Of these, $800$ contain clarified implicit references, while $73$ do not require clarification. Within the $800$ opinions, there are a total of $1113$ written clarified implicit references.
@@ Appendix E LLM Sources
Appendix F Human Evaluation Instruction
You will be given a pair of summaries (a Background Summary and an Opinion Summary).
Please evaluate the overall quality of this pair of summaries based on the following three criteria. Provide a single, holistic score from 1 to 5 using the Likert scale (5 = Best, 1 = Worst).
1. Background Summary: Does the summary provide the necessary context to understand the discussion?
2. Opinion Summary: Does the summary clearly provide the participants’ viewpoints with implicit references successfully clarified?
3. Overall Performance: Do the two summaries work together effectively for the readers?
Our human evaluation is conducted by the same PhD candidates who constructed the benchmark. We choose these experts as they possess a deep and consistent understanding of the task’s goals, which is already demonstrated by their high agreement during the benchmark creation phase. This evaluation process yields a Cohen’s Kappa coefficient of 0.74, indicating substantial inter-annotator agreement.
Appendix G Prompts and Instructions
G.1 Unified Parameter Settings
In this work, all 12 evaluated LLMs use consistent parameter settings for all prompts and instructions: max_token=4096 and temperature=0. No other default parameters are modified.
G.2 Atomic Fact Decomposition Instruction
@@ G.3 Structured Prompts
G.4 Self-Reflection Instructions
@@ G.5 Verification Prompts
G.6 Error Detection Instruction